OpenStudy (anonymous):

Find the coefficient of x^3 in the expansion of (1-3x)(1+2x)^6 Please Help! Thank you!

OpenStudy (blockcolder):

\[(1+2x)^6=\sum_{i=0}^6\binom{6}{i}(2x)^i\\ (1-3x)(1+2x)^6=(1-3x)\left[\sum_{i=0}^6\binom{6}{i}2^ix^i\right]=\sum_{i=0}^6\binom{6}{i}2^ix^i-\sum_{k=0}^63\binom{6}{k}2^kx^{k+1}\] To get an x^3, you need i=3 and k=2.

OpenStudy (blockcolder):

Thus, the coefficient of x^3 is \[\binom{6}{3}2^3-3\binom{6}{2}2^2\]